OpenStudy (anonymous):

Rate of change question... Let f(x)= (1/x) a) Find the average rates of change of f(x) over the intervals [1,1.1], [1,1.01], and [1,1.001]. Thanks

OpenStudy (freckles):

To find the average rate of change over the interval [a,b] you simply do: \[\frac{f(a)-f(b)}{a-b}\]

OpenStudy (anonymous):

I know the formula, I just am having a hard time solving it because the X values are all the same

OpenStudy (anonymous):

No I read that wrong initially. You are given intervals \(x \in [1, 1.1]\), so you need to find the corresponding f(x) values when \(x=1\) and when \(x=1.1\) \(f(1)=\dfrac{1}{1}\) \(f(1.1)=\dfrac{1}{1.1}\)

OpenStudy (anonymous):

do the same for the other 2 intervals

OpenStudy (anonymous):

thank you